# Synthesizing Grid Data with Cyber Resilience and Privacy Guarantees

### v1

@@ V-A Substantiating Attacks Calibrated on DP Data
Table I collects the damage of load redistribution attacks on the actual $\mathbf{d}$ and synthetic $\tilde{\mathbf{d}}$ loads. The synthetic parameters are generated using the standard post-processing steps in (4) with no cyber resilience guarantee. The results reveal that the load redistribution attacks are as effective on synthetic parameters as on the original loads, thus motivating the cyber resilient obfuscation by means of Alg. 1 and 2.
@@ V-B Insights from the Small PJM 5-Bus Testbed
TABLE II: OPF costs induced on synthetic load vectors $\tilde{\mathbf{d}}_{\text{cro}}$ for varying trade-off parameters $\beta$ and attack assumptions. Attack magnitudes $\eta=5\%$.
@@ V-C Large-Scale Applications with CRO-Exp
The post-processing optimization (1) in CRO is difficult to scale to large systems. As shown in Fig. 2, the number of variables and complementarity constraints grow with the size of the testbed. The CRO-Exp Alg. 2 reduces the problem by several orders of magnitude as it only considers a subset of $\tau$ worst-case constraints in the attack. Fig. 3 shows the damage of attacks calibrated on synthetic loads released by CRO-Exp for three large testbeds. The increase of $\tau$ reduces the attack damage. Notably, $\tau=5$ suffices to minimize the attack damage, showing no improvement of cyber resilience beyond this threshold. This is due to the fact that only the attacks on a limited number of constraints can greatly increase the OPF cost. Moreover, the selection of the worst-case constraints in Step 3 of Alg. 2 becomes less informative with more noise, which only increases in $\tau$, as per Theorem 5.
Figure 2: Number of variables and complementarity constraints in CRO and CRO-Exp ($\tau=5$) post-processing across four test systems (log-scale).
Figure 3: Outcomes of the BO load redistribution attack calibrated on synthetic CRO-Exp loads for varying number of the worst-case constraints $\tau$. $\tau=0$ means the synthetic dataset generated by the standard post-processing in (4). Adjacency $\alpha$ are determined in percentages of the average load in the testbed. Attack magnitudes are for $\eta=15\%$ in IEEE 14-bus system and $\eta=5\%$ in IEEE 24-bus and 118-bus systems. Red lines represent the mean value, the blue area represents the 80% confidence interval.
@@ VI Conclusion
For any feasible load vectors $\mathbf{d}$, we have
$C_{\text{opf}}^{\text{BO}}(\mathbf{d})$ $=\underset{{\mathbf{x}}}{\text{minimize}}\quad\mathbf{c}^{\top}\mathbf{x}$ (12)
subject to $\mathbf{a}_{k}^{\top}\mathbf{x}+\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\gamma}^{\star})+e_{k}\leqslant 0\;:\omega_{k}$
$C_{\text{opf}}^{\text{RO}}(\mathbf{d})$ $=\underset{{\mathbf{x}}}{\text{minimize}}\quad\mathbf{c}^{\top}\mathbf{x}$ (13)
subject to $\mathbf{a}_{k}^{\top}\mathbf{p}+\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\delta}_{k}^{\star})+e_{k}\leqslant 0\;:\xi_{k},$
where ${\gamma}^{\star}$ and $\{\bm{\delta}_{1}^{\star},...,\bm{\delta}_{k}^{\star}\}$ are the optimal attack vectors, belonging to the same admissible set $\Delta$. We show the correspondence between the objectives via Lagrange duality of problems (12) and (13). Their dual problems are
$\underset{{\omega_{k}}}{\text{maximize}}\quad$ $\textstyle\sum_{k}\omega_{k}(\mathbf{b}_{k}^{\top}\bm{\gamma}^{\star}+\mathbf{b}_{k}^{\top}\mathbf{d}+e_{k})$ (14a)
subject to $\mathbf{c}+\textstyle\sum_{k}\omega_{k}\mathbf{a}_{k}=\bm{0}\quad\forall k,$ (14b)
$\underset{{\xi_{k}}}{\text{maximize}}\quad$ $\textstyle\sum_{k}\xi_{k}(\mathbf{b}_{k}^{\top}\bm{\delta}_{k}^{\star}+\mathbf{b}_{k}^{\top}\mathbf{d}+e_{k})$ (15a)
subject to $\mathbf{c}+\textstyle\sum_{k}\xi_{k}\mathbf{a}_{k}=\bm{0}\quad\forall k,$ (15b)
respectively. To relate the costs, we thus need to show that
$\textstyle\sum_{k}\xi_{k}(\mathbf{b}_{k}^{\top}\bm{\delta}_{k}^{\star}+\mathbf{b}_{k}^{\top}\mathbf{d}+e_{k})$
$\quad\quad\quad\geqslant\textstyle\sum_{k}\omega_{k}(\mathbf{b}_{k}^{\top}\bm{\gamma}^{\star}+\mathbf{b}_{k}^{\top}\mathbf{d}+e_{k}).$ (16)
From dual constraints (15b) and (14b), we have the equality
$\textstyle\sum_{k}\xi_{k}\mathbf{a}_{k}=\textstyle\sum_{k}\omega_{k}\mathbf{a}_{k},\quad\forall k=1,\dots,K,$ (17)
for which $\xi_{k}=\omega_{k},\forall k=1,\dots,K$ is one possible solution. By Assumption 1, the OPF problem (1) satisfies LICQ, and so do problems (12) and (13) for any attack vector. Per [19, Theorem 2], for any feasible dispatch $\mathbf{p}$, the dual solution to problems (12) and (13) is unique. Hence, $\xi_{k}=\omega_{k}$ is not only one but also unique solution to (17). Since the duals are non-negative in optimality, it suffices to show that
$\mathbf{b}_{k}^{\top}\bm{\delta}_{k}^{\star}\geqslant\mathbf{b}_{k}^{\top}\bm{\gamma}^{\star},\quad\forall k=1,\dots,K,$ (18)
for equation (16) to hold. We recall that the optimal attack vectors $\bm{\delta}_{1}^{\star},\dots,\bm{\delta}_{K}^{\star}$ come from the robust optimization (7), so the left-hand side is $\mathbf{b}_{k}^{\top}\bm{\delta}_{k}^{\star}=\underset{{\bm{\delta}_{k}\in\Delta}}{\text{maximize}}\;\mathbf{b}_{k}^{\top}\bm{\delta}_{k}$. On the other hand, the worst case for this inequality is when $\bm{\gamma}$ maximizes the right-hand side of (18), so we need to show that
$\underset{{\bm{\delta}_{k}\in\Delta}}{\text{maximize}}\;\mathbf{b}_{k}^{\top}\bm{\delta}_{k}\geqslant\underset{{\bm{\gamma}\in\Delta}}{\text{maximize}}\;\mathbf{b}_{k}^{\top}\bm{\gamma},\quad\forall k=1,\dots,K.$ (19)
Since the attack vectors belong to the same set $\Delta$, (19) holds as equality in the worst case. Therefore, inequality (18), and hence (16), always hold, completing the proof.

### v2

@@ V-A Substantiating Attacks Calibrated on DP Data
Table I collects the damage of load redistribution attacks. The synthetic loads are generated using the standard post-processing (4) with no cyber resilience guarantee. The results reveal that the load redistribution attacks are as effective on synthetic loads as on the original loads, motivating the cyber resilient obfuscation by means of Alg. 1 and 2.
@@ V-B Insights from the Small PJM 5-Bus Testbed
TABLE II: OPF costs induced on synthetic load vectors $\tilde{\mathbf{d}}_{\text{cro}}$ for varying trade-off parameters $\beta$ and load adjacency $\alpha$. Attack magnitude $\eta=5\%$.
@@ V-C Large-Scale Applications with CRO-Exp
The post-processing optimization (1) in CRO is difficult to scale to large systems. As shown in Fig. 2, the number of variables and complementarity constraints grow with the size of the testbed. The CRO-Exp Alg. 2 reduces the problem by at least one order of magnitude to a similar level as the standard post-processing, since it only considers a subset of $\tau$ worst-case constraints in the attack. Fig. 3 shows the damage of attacks calibrated on synthetic loads released by CRO-Exp for three large testbeds. The increase of $\tau$ reduces the attack damage. Notably, $\tau=5$ suffices to minimize the attack damage, showing no improvement of cyber resilience beyond this threshold. This is due to the fact that only the attacks on a limited number of constraints can greatly increase the OPF cost. Moreover, the selection of the worst-case constraints in Step 3 of Alg. 2 becomes less informative with more noise, which only increases in $\tau$, as per Theorem 5.
Figure 2: Num. of variables and complementarity constraints in CRO, CRO-Exp ($\tau$=5) and standard post-processing (4) across four testbeds (log-scale).
Figure 3: Outcomes of the BO load redistribution attack calibrated on synthetic CRO-Exp loads for varying number of the worst-case constraints $\tau$. The damage in percentage is computed as $(C_{\text{att}}^{\text{BO}}(\tilde{\mathbf{d}})-C_{\text{opf}}(\tilde{\mathbf{d}}))/C_{\text{opf}}(\tilde{\mathbf{d}})\times 100$. $\tau=0$ means the synthetic dataset generated by the standard post-processing in (4). Adjacency $\alpha$ are determined in percentages of the average load in the testbed. Attack magnitudes are $\eta=15\%$ in IEEE 14-bus system and $\eta=5\%$ in IEEE 24-bus and 118-bus systems. Red lines represent the mean value, the blue area represents the 80% confidence interval.
@@ VI Conclusion
Consider two perturbed OPF problems formulated on the same vector $\mathbf{d}$, one resulting from the BO attack (5)
$C_{\text{opf}}^{\text{BO}}(\mathbf{d})$ $=\underset{{\mathbf{x}}}{\text{minimize}}\quad\mathbf{c}^{\top}\mathbf{x}$ (12a)
subject to $\mathbf{a}_{k}^{\top}\mathbf{x}\leqslant-\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\gamma}^{\star})-e_{k},\;\forall k$ (12b)
and one from the RO approximation of the attack in (7)
$C_{\text{opf}}^{\text{RO}}(\mathbf{d})$ $=\underset{{\mathbf{x}}}{\text{minimize}}\quad\mathbf{c}^{\top}\mathbf{x}$ (13a)
subject to $\mathbf{a}_{k}^{\top}\mathbf{x}\leqslant-\mathbf{b}_{k}^{\top}(\mathbf{d}+\bm{\delta}_{k}^{\star})-e_{k},\;\forall k$ (13b)
with perturbations $\bm{\gamma}^{\star},\bm{\delta}_{1}^{\star},\dots,\bm{\delta}_{K}^{\star}\in\Delta.$
To show that the optimal value of (12a) is upper-bounded by the optimal value of (13a), we need to establish that the feasible set (13b) is a subset of (12b). This is per the global inequality in perturbation analysis of convex programs [19, §5.6, Eq. (5.57)]. Inspecting (12b) and (13b), observe that this is the case when
$\mathbf{b}_{k}^{\top}\bm{\delta}_{k}^{\star}\geqslant\mathbf{b}_{k}^{\top}\bm{\gamma}^{\star},\quad\forall k=1,\dots,K.$ (14)
The attack vectors $\bm{\delta}_{1}^{\star},\dots,\bm{\delta}_{K}^{\star}$ come from the RO, so the left-hand side of (14) is given by the following optimization:
$\mathbf{b}_{k}^{\top}\bm{\delta}_{k}^{\star}=\max_{\bm{\delta}_{k}\in\Delta}\;\mathbf{b}_{k}^{\top}\bm{\delta}_{k},\quad\forall k=1,\dots,K.$ (15)
At the same time, the right-hand side of (14) can be represented by the following optimization problem:
$\mathbf{b}_{k}^{\top}\bm{\gamma}^{\star}=\max_{\bm{\gamma}_{k}\in\Delta}\;$ $\mathbf{b}_{k}^{\top}\bm{\gamma}_{k}$ (16a)
$\quad\quad\quad\quad\quad\forall k=1,\dots,K$
s.t. $\bm{\gamma}_{k}=\bm{\gamma}^{\star}$ (16b)
Although trivial, this optimization problem allows us to clearly relate both sides of inequality (14) by relating problems (15) and (16). They are similar except for the additional consensus constraint (16b). Since $\bm{\gamma}^{\star}\in\Delta$ by design, the feasible set of $\bm{\gamma}_{k}$ is the subset of that for $\bm{\delta}_{k}$. Hence, we can conclude that the optimal value of (15) is greater or equal than that of (16a). Therefore, inequality (14) holds and (13b) is indeed a subset of (12b), completing the proof.